# Improvement of Pólya’s conjecture for balls and cylinders

### v1

@@ Abstract.
Laplacian, eigenvalues, Weyl’s law, Pólya’s conjecture
@@ 1. Introduction
This improves upon the remainder estimate $o(\mu^{d-1})$ in the Weyl conjecture for general domains and manifolds with boundary (see, for example, Ivrii [8]). Thus, the asymptotic behavior of $\mathscr{N}_{\mathcal{C}_{R,L}^{d}}(\mu)$ as $\mu\to\infty$ is quite clear.
@@
In 1961, Pólya [13] proved his conjecture (1.2) for planar tiling domains. While his proof for the Neumann case relied on an additional “regular tiling” assumption, this restriction was removed by Kellner [10] in 1966. In 1984, Urakawa [14] generalized Pólya’s result to bounded domains in $\mathbb{R}^{d}$ for the Dirichlet case. This derivation, which incorporated an extra factor of the lattice packing density, confirmed the conjecture for higher-dimensional tiling domains. In 1997, Laptev [11] showed that the Dirichlet Pólya’s conjecture holds for the product $\Omega_{1}\times\Omega_{2}$ whenever it holds for $\Omega_{1}\subset\mathbb{R}^{d_{1}}$ with $d_{1}\geq 2$, and $\Omega_{2}\subset\mathbb{R}^{d_{2}}$ is of finite Lebesgue measure. In 2023, Freitas and Salavessa [5] explored Laptev’s approach and showed the existence of classes of non-tiling domains satisfying Pólya’s conjecture in any dimension, in both the Euclidean and non-Euclidean cases (see also Freitas, Lagacé, and Payette [4]). Recently, He and Wang [7] proved Pólya’s conjecture for certain thin products of Euclidean domains of the form $(a\Omega_{1})\times\Omega_{2}$ for small $a>0$. We refer interested readers to [4, 7] and the references therein for further results on unions and products of domains. Recent progress has also been made on variants of the conjecture. Notably, Filonov [1] tackled the Neumann case for planar convex domains, and Jiang and Lin [9] obtained an $\epsilon$-loss version of the Dirichlet case for bounded Lipschitz domains in $\mathbb{R}^{d}$.
In a remarkable 2023 paper, Filonov, Levitin, Polterovich, and Sher [2] achieved a breakthrough by proving Pólya’s conjecture for the disk. This work resolved a long-standing open problem, establishing the disk as the first non-tiling planar domain for which the celebrated conjecture is known to hold. Their work also encompassed the confirmation of the conjecture for arbitrary planar sectors and, in the Dirichlet case, for balls of any dimension, showcasing the power and versatility of their methodology. Their very recent follow-up work [3] has further expanded the frontier by verifying the Dirichlet Pólya’s conjecture for annuli, another fundamental and geometrically non-trivial domain.
@@
In another word, we are curious about the “in-between” case (that is, between “large $\mu$” and “nonnegative $\mu$”). We seek to characterize the behavior for values of $\mu$ that are positive yet not necessarily large, thereby addressing the gap between these two previously understood cases.
Very recently, Filonov, Levitin, Polterovich and Sher’s [3, Theorem 1.5] provides such an improvement for the Dirichlet Laplacian on the unit disk:
@@
For $\mu\geq 1/\omega_{0}$, this result establishes a bound that is strictly sharper than that of Pólya’s conjecture.
@@
In Section 3.2 we prove an improvement of Pólya’s conjecture for product domains slightly more general than cylinders. Theorem 1.9 follows directly from Theorems 3.2 and 3.5 below combined with [2, Theorem 1.2 and Corollary 1.6].
This result directly confirms the Neumann Pólya’s conjecture for cylinders in $\mathbb{R}^{3}$ (while the Dirichlet Pólya’s conjecture is already known to hold for cylinders in $\mathbb{R}^{d}$, $d\geq 3$, as a consequence of [2, Theorem 1.2] and [11, Theorem 2.8]).
@@ Acknowledgments
C. Miao is supported by the National Key R&D Program of China (no. 2022YFA1005700) and the NSFC (no. 12371095). J. Guo is supported by the NSFC (no. 12571110 and 12341102).

### v2

@@ Abstract.
Laplacian, eigenvalues, Weyl’s law, Pólya’s conjecture, improvement
@@ 1. Introduction
This improves upon the remainder estimate $o(\mu^{d-1})$ in the Weyl conjecture for general domains and manifolds with boundary (see, for example, Ivrii [8]). Thus, we now have a better understanding of the behavior of $\mathscr{N}_{\mathcal{C}_{R,L}^{d}}(\mu)$ as $\mu\to\infty$.
@@
In 1961, Pólya [13] proved his conjecture (1.2) for planar tiling domains. While his proof for the Neumann case relied on an additional “regular tiling” assumption, this restriction was removed by Kellner [10] in 1966. In 1984, Urakawa [14] generalized Pólya’s result to bounded domains in $\mathbb{R}^{d}$ for the Dirichlet case. This derivation, which incorporated an extra factor of the lattice packing density, confirmed the conjecture for higher-dimensional tiling domains. In 1997, Laptev [11] showed that the Dirichlet Pólya’s conjecture holds for the product $\Omega_{1}\times\Omega_{2}$ whenever it holds for $\Omega_{1}\subset\mathbb{R}^{d_{1}}$ with $d_{1}\geq 2$, and $\Omega_{2}\subset\mathbb{R}^{d_{2}}$ is of finite Lebesgue measure. In 2023, Freitas and Salavessa [5] explored Laptev’s approach and showed the existence of classes of non-tiling domains satisfying Pólya’s conjecture in any dimension, in both the Euclidean and non-Euclidean cases (see also Freitas, Lagacé, and Payette [4]). Recently, He and Wang [7] proved Pólya’s conjecture for certain thin products of Euclidean domains of the form $(a\Omega_{1})\times\Omega_{2}$ for small $a>0$. We refer interested readers to [4, 7] and the references therein for further results on unions and products of domains. Recent progress has also been made on variants of the conjecture. Notably, Filonov [1] tackled the Neumann case for planar convex domains, and Jiang and Lin [9] proved an $\epsilon$-loss version of the Dirichlet case for bounded Lipschitz domains in $\mathbb{R}^{d}$ based on a uniform estimate of the remainder of Weyl’s law, and provided in all dimensions a class of general domains satisfying the Dirichlet Pólya’s conjecture.
In a remarkable 2023 paper, Filonov, Levitin, Polterovich, and Sher [2] achieved a breakthrough by proving Pólya’s conjecture for the disk. This work resolved a long-standing open problem, establishing the disk as the first non-tiling planar domain for which the celebrated conjecture is known to hold. Their work also encompassed the confirmation of the conjecture for arbitrary planar sectors and, in the Dirichlet case, for balls of any dimension, showcasing the power and versatility of their methodology. Their very recent follow-up work [3] has further expanded the frontier by verifying the Dirichlet Pólya’s conjecture for annuli, another fundamental and geometrically non-trivial domain, thus providing the first verification of the conjecture for a non-simply connected planar domain.
@@
In another word, we are curious about the “in-between” case (that is, between “large $\mu$” and “nonnegative $\mu$”). We seek to characterize the behavior for values of $\mu$ that are positive yet not necessarily large, thereby addressing the gap between these two previously studied cases.
Very recently, Filonov, Levitin, Polterovich and Sher’s [3, Theorem 1.5] provided such an improvement for the Dirichlet Laplacian on the unit disk:
@@
For $\mu\geq 1/\omega_{0}$, this result establishes a bound that is strictly sharper than that of Pólya’s conjecture. This improvement is based on a refined lattice-point counting estimate, which is a pivot part of their proof of the Dirichlet Pólya’s conjecture for planar annuli.
Later, Jiang and Lin’s [9, Theorem 2.3] obtained an improvement for the Dirichlet Laplacian on the product domain $\Omega_{1}\times\Omega_{2}\subset\mathbb{R}^{d_{1}}\times\mathbb{R}$, with $d_{1}\geq 1$ and $\Omega_{1}$, $\Omega_{2}$ bounded, under the assumption that the Dirichlet Pólya’s conjecture holds for $\Omega_{1}$.
@@
In Section 3.2 we prove an improvement of Pólya’s conjecture for product domains slightly more general than cylinders, though we do not attempt to extend the result to a fully general setting. Theorem 1.9 follows directly from Theorems 3.2 and 3.5 below combined with [2, Theorem 1.2 and Corollary 1.6].
While our Theorem 3.2 addresses the Dirichlet case in a manner analogous to Jiang and Lin’s Theorem 2.3, it relies only on elementary analysis and leads to distinct bounds. The corresponding Neumann case is handled in our Theorem 3.5.
From the proof of Theorem 3.5, it is easy to observe that: the Neumann Pólya’s conjecture holds for the product $\mathscr{D}\times I$ whenever it holds for $\mathscr{D}\subset\mathbb{R}^{d-1}$, with $d\geq 3$, and $I$ is a finite interval.
Theorem 1.9 directly confirms the Neumann Pólya’s conjecture for cylinders in $\mathbb{R}^{3}$ (while the Dirichlet Pólya’s conjecture is already known to hold for cylinders in $\mathbb{R}^{d}$, $d\geq 3$, as a consequence of [2, Theorem 1.2] and [11, Theorem 2.8]).
@@ Acknowledgments
We thank Prof. Renjin Jiang for helpful comments, which included drawing our attention to their result [9, Theorem 2.3]. C. Miao is partially supported by the National Key R&D Program of China (no. 2022YFA1005700) and the NSFC (no. 12371095). J. Guo is partially supported by the NSFC (nos. 12571110 and 12341102).